3/(4x+3) = 1/4x ????????????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

LHS =3/(4x+3) ≠ 1/4x
The correct statement is 3/(4x+3) = 3/(4x+3)

Thanks